OpenStudy (anonymous):

If g(x) = 3(x - 2), find the value of x if g(x) = 6. A. 12 B. 2.6 C. 8 D. 4

OpenStudy (anonymous):

Substitute the values. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there?

OpenStudy (jacob902):

true

OpenStudy (anonymous):

yea thnx :)

Nnesha (nnesha):

g(x) equal 6 not just x

OpenStudy (anonymous):

yw :)

Nnesha (nnesha):

that's totally different question

Nnesha (nnesha):

\[g(x) \cancel= 4\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @iloveyou;* If g(x) = 3(x - 2), find the value of x if g(x) = 6. A. 12 B. 2.6 C. 8 D. 4 \(\color{blue}{\text{End of Quote}}\) for this one replace g(x) by 6 distribute parentheses by 3

OpenStudy (anonymous):

thnx :)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Rachel_K Substitute the values. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there? \(\color{blue}{\text{End of Quote}}\) \[4 \cancel = 4^2 +3\]

OpenStudy (anonymous):

g(4)=4^2+3=16+3=19